OpenStudy (anonymous):

5. Which equation models the problem below if c represents the number of babies whose eyes are closed? There are 12 babies in a photograph, and 9 of them have their eyes open. How many babies have their eyes closed? (Points : 1) 12 + c = 9 9 + c = 12 c – 9 = 12 9 – c = 12

OpenStudy (anonymous):

c-9=12 9-c=12 ^at the ending

OpenStudy (micahm):

so what are you looking for

OpenStudy (anonymous):

How many babys eyes are either closed, or opened. i believe

OpenStudy (anonymous):

Never mind, but thank you for your consideration of helping, I appreciate it.

OpenStudy (micahm):

well have you look at A or D is what i choose